# Tight Algorithm for Connected Odd Cycle Transversal Parameterized by Clique-width

### 12 Feb 2024 v1

@@ 5.4 Isolation lemma
Assume that $G$ contains a connected odd cycle transversal of size $\overline{b}$. Fix $W=8\cdot|\mathcal{V}^{CJ}|$, and assume that $\omega$ is chosen uniformly at random in $[W]^{\big{(}\mathcal{V}^{CJ}\times[4]\big{)}}$. Let $\tau$ be a minimum weight valid action-sequence of cost $\overline{b}$ generating the pair $([0],c)$ for any coloring $c$. Then $\tau$ is unique with probability at least one half.
@@
Let $P$ be the family of all valid action-sequences of cost $\overline{b}$ generating the pair $([0],c)$ for some coloring $c$. The weight of each such sequence is given by $\sum\limits_{x\in\mathcal{V}^{CJ}}\omega(x,\tau(x))$. Let $U=\mathcal{V}^{CJ}\times[4]$, and for $\tau\in P$, let $F_{\tau}=\{(x,\tau(x))\colon x\in\mathcal{V}^{CJ}\}$. Let =$\mathcal{F}=\{F_{\tau}\colon\tau\in P\}$. Then $\tau\mapsto F_{\tau}$ is a bijection from $P$ to $\mathcal{F}$. Moreover, it holds that $\omega(\tau)=\omega(F_{\tau})$ for all $\tau\in P$. Finally, by Lemma 5.23, it holds that $\omega$ isolates $\mathcal{F}$ with probability
@@ 5.5 Correctness of the algorithm
Let $W=8\cdot|\mathcal{V}^{CJ}|$. First, choose a weight function $\omega:\mathcal{V}^{CJ}\times[4]\rightarrow W$, by choosing $\omega(x,\ell)\in[W]$ independently and uniformly at random for each $(x,\ell)\in\mathcal{V}^{CJ}\times[4]$. For each vertex $v\in V$, the algorithm sets $v_{0}=v$ and computes all tables $T{\big{[}{x,{b},{w}}\big{]}}$ for all $x\in\mathcal{V}$, and all values ${b}\in[\overline{b}]$ and ${w}\in[4\cdot|\mathcal{V}^{CJ}|\cdot W]$ in a bottom-up manner over $\mathcal{T}$. The algorithm returns yes, if for any choice of $v_{0}$ there exists a coloring $c\in\mathscr{C}$ and a weight ${w}$ such that $T{\big{[}{r,\overline{b},{w}}\big{]}}\big{[}{([0],c)}\big{]}=1$. Otherwise, it returns no.
@@
On the other hand, if a solution of size $\overline{b}$ exists. Let $v_{0}$ be any vertex in any solution of size $\overline{b}$. Then by Lemma 5.24 there exists a weight ${w}$ and a coloring $c$, such that with probability at least $1/2$ there exists a unique action-sequence at $r$ of cost $\overline{b}$ and weight ${w}$ generating the pair $([0],c)$, which implies by Corollary 5.21 that $T{\big{[}{r,\overline{b},{w}}\big{]}}\big{[}{[0],c}\big{]}=1$. Hence, with probability at least $1/2$ the algorithm outputs yes. ∎
@@ 6.1 Construction of the graph
By adding the triangle $\Delta({u},{v})$ for two vertices $u$ and $v$, we mean that we add the edge $\{u,v\}$, and we add a new vertex $w_{u,v}$ and make it adjacent to $u$ and $v$ only. When we say we add a triangle at a vertex $v$, we mean that we add two new vertices $u_{v},w_{v}$, and make them adjacent to each other and to $v$ only. Vertices created when adding triangles are called triangle vertices. For a set of vertices $S\subseteq V$, we denote by $\Delta{S}$ the set consisting of $S$ and all vertices added to create triangles at single vertices of $S$ or between pairs of vertices in $S$.
We build $G$ by combining components of constant size, called gadgets. We distinguish three kinds of gadgets in $G$, namely path gadgets, decoding gadgets, and clause gadgets. We distinguish a single vertex $r$ in $G$ called the root vertex, with a triangle added to it, and another single vertex $r_{\boldsymbol{B}}$. When we say that we color a vertex white in $G$, it means that we add an edge between this vertex and $r_{\boldsymbol{B}}$. When we say we color a vertex $u$ black, we mean that we create a new vertex $u^{\prime}$, we color it white, and we add an edge between $u$ and $u^{\prime}$. We also distinguish two vertices $g_{1}$ and $g_{2}$ in $G$, called guards. We add a triangle at each of them.
@@
Let $t=\left\lceil t_{0}\cdot\log_{12}(2)\right\rceil$. Let $n^{\prime}=s\cdot t$. The graph $G$ consists of $n^{\prime}$ path-like structures, called path sequences. Each path sequence is a sequence of $c=(11\cdot n^{\prime}+1)\cdot m$ path gadget. Consecutive path gadgets are connected by cuts of size $9$, given as bicliques between three vertices of one gadget, and three of the other, giving a path sequence its narrow path-like structure. The first path gadget of each path sequence is adjacent to $g_{1}$, while the last path gadget is adjacent to $g_{2}$.
Let $P_{1},\dots,P_{n^{\prime}}$ be the set of all path sequences in $G$. We group them into $s$ groups of size $t$, calling each group a path bundle. Hence, the $i$th path bundle $\mathcal{B}_{i}$ consists of the path sequences $P_{(i-1)\cdot t+1}\dots P_{i\cdot t}$, and it corresponds to the variable group $V_{i}$. For each $i\in[s],j\in[c]$, we call the set of all $j$th path gadgets on all path sequences in the path bundle $\mathscr{B}_{i}$ a bundle $B_{i}^{j}$.
For a path sequence $P_{i}$, let $X_{i}^{1},\dots X_{i}^{c}$ be the path gadgets on the path sequence $P_{i}$ from left to right. We divide each path sequence $P_{i}$ into $(11\cdot n^{\prime}+1)$ groups of size $m$ each, called segments $S_{i}^{1},\dots,S_{i}^{11\cdot n^{\prime}+1}$, where $S_{i}^{j}=X_{i}^{(j-1)\cdot m+1},\dots,X_{i}^{j\cdot m}$. For $j\in[11\cdot n^{\prime}+1]$ We call the set of the $j$th segment over all path sequences a section $\mathcal{S}_{j}$, i.e. $\mathcal{S}_{j}=\{S_{i}^{j}\colon i\in[n^{\prime}]\}$. Finally, for $j\in[c]$, we call the set of the $j$-th path gadget on each path sequence a column $\operatorname{Col}_{j}=\{X_{i}^{j}\colon i\in[n^{\prime}]\}$. See Figure 1 for illustration.
@@
We attach a decoding gadget $D_{i}^{j}$ to each bundle $B_{i}^{j}$ for $i\in[s]$ and $j\in[c]$, and we attach a clause gadget $Z_{j}$ to each column $\operatorname{Col}_{j}$ for $j\in[c]$. The decoding gadgets are only adjacent (possibly through triangles) to vertices in their corresponding path gadget, to the clause gadget in their column, and to $r$. A clause gadget is only adjacent to vertices in decoding gadgets attached to bundles in its column.
Let =$\mathscr{I}=\{\boldsymbol{C},\boldsymbol{D},\boldsymbol{B},\boldsymbol{W}\}$ be a set of states. We define the set of tuples $\mathcal{S}=\{s_{1},\dots s_{12}\}\subseteq\mathscr{I}^{3}$ given by the following list:
@@
$\big{|}\mathcal{S}^{B}\big{|}=12^{t}\geq 12^{t_{0}/\log 12}\geq 2^{t_{0}}\geq\big{|}\{\mathbf{T},\mathbf{F}\}^{V_{i}}\big{|},$
we can fix an arbitrary injective mapping $\tau:\{\mathbf{T},\mathbf{F}\}^{V_{i}}\rightarrow\mathcal{S}^{[t]}$. Since both the set of the variables, and all path gadgets in each bundle are ordered (the former by index, and the latter by the index of the path sequence they belong to), we can assume that this mapping is defined over all variable groups $V_{i}$. Now we describe the different gadgets.
@@
A path gadget $X$ consists of six simple path gadgets $Y_{1},Y_{2},Y_{3},Y_{4},Y_{5},Y_{6}$. We denote a vertex $x$ in $Y_{i}$ by $x^{i}$. Hence, by $v^{1}$ we denote the vertex $v$ in $Y_{1}$, and by $x_{\boldsymbol{C}}^{3}$ we denote the vertex $x_{\boldsymbol{C}}$ in $Y_{3}$. We call the vertices $v^{1},v^{2},v^{3}$ the entry vertices, and the vertices $v^{4},v^{5},v^{6}$ the exit vertices of $X$. These vertices play a special role in $G$, since they connect consecutive path gadgets on a path sequence. In particular, we add a biclique between the exit vertices of each path gadget (except for the last one), and the entry vertices of the following path gadget. The entry vertices of the first path gadget of each path sequence are all adjacent to $g_{1}$, and the exist vertices of the last path gadget are all adjacent to $g_{2}$.
For each $s\in\mathscr{I}$, we add a vertex $z_{s}$ to $X$ and make it adjacent to $r$. We call the resulting vertices the transition vertices of $X$. We add all triangles $\Delta({z_{s}},{z_{s^{\prime}}})$ for all $s\neq s^{\prime}\in\mathcal{S}$. Finally, for each $s\in\mathcal{S}$, let $s=(\chi_{1},\chi_{2},\chi_{3})$ and $\operatorname{next}(s)=(\chi_{4},\chi_{5},\chi_{6})$, where $\chi_{i}\in\mathscr{I}$ for all $i\in[6]$. For each $i\in[6]$ we add all triangles $\Delta({z_{s}},{x_{\chi}^{i}})$, for all $\chi\neq\chi_{i}\in\mathscr{I}$. That means we add a triangle between a transition vertex $z_{s}$ and a state vertex $x_{\chi}^{i}$ in $Y_{i}$, if $\chi_{i}\neq\chi$. This completes the description of a path gadget. Given a vertex $v$ in a path gadget $X$, if not clear from the context, we write $v(X)$ to indicate the path gadget it belongs to. In total, a path gadget contains $6\cdot 10+12=72$ non-triangle vertices.
@@
Let $D=D_{i}^{j}$, and let $X_{1},\dots X_{t}$ be the path gadgets on the bundle $B_{i}^{j}$. For each mapping $\pi\in\Pi$ and for each $i\in[t]$, and for $s\in\mathcal{S}$, we add the triangle $\Delta({z_{s}(X_{i})},{u_{\pi}})$ if it holds that $\pi(i)\neq s$. Given a vertex $v$ in a decoding gadget $D$, if not clear from the context, we write $v(D)$ to indicate the decoding gadget it belongs to. In total, a decoding gadget contains $2\cdot 12^{t}$ non-triangle vertices.
@@
Now we define the target value ${b}$ in the resulting instance $(G,{b})$ of the Connected Odd Cycle Transversal problem. First, we define a family of disjoint components $\mathcal{F}$ in $G$ (mostly the gadgets of $G$). For each component $X\in\mathcal{F}$, we will define a family of disjoint subsets of $X$ (denoted $\beta(X)$). For each set $S\in\beta(X)$, we will define a lower bound ${b}(S)$ on the size of a connected odd cycle transversal in $G[S]$. We define ${b}(X)$ as the sum of ${b}(S)$ for all $S\in\beta(X)$. Finally, we define ${b}$ as the sum of ${b}(X)$ for all $X\in\mathcal{F}$
@@ 6.2 Proof of correctness
Now we are ready to describe the linear $k$-expression of $G$. First, we introduce the vertices $r,r_{\boldsymbol{B}},g_{1}$ and $g_{2}$ with labels $\ell_{\operatorname{root}}=n^{\prime}+3$, $\ell_{\operatorname{black}}=n^{\prime}+4$, $\ell_{\operatorname{g_{1}}}=n^{\prime}+5$, and $\ell_{\operatorname{g_{2}}}=n^{\prime}+6$ respectively. These are the only vertices that will be assigned the corresponding labels at any timestamp, and will never be relabeled. Hence, for a vertex $u$ of a unique label at any timestamp, we can make it adjacent to any of them through a join operation without changing the adjacency of any other vertex of the graph. We introduce the private neighbors of $r,g_{1},$ and $g_{2}$, using a different label for each created vertex. We create the corresponding triangles using join operations, and then we forget all these private neighbors. Finally, we reserve the label $\ell_{\operatorname{white}}=n^{\prime}+7$ for creating white vertices when coloring some vertices black. For a vertex $v$ of a unique label $\ell$, when we say we color $v$ white we mean that we join the labels $\ell$ and $\ell_{\operatorname{black}}$, while by coloring $v$ black, we mean that we create a vertex $w$, labeled $\ell_{\operatorname{white}}$, and then apply two join operations: one between $\ell_{\operatorname{white}}$ and $\ell_{\operatorname{black}}$, and another between $\ell_{\operatorname{white}}$ and $\ell$. After that we forget the label $\ell_{\operatorname{white}}$.
We divide the rest of the construction into phases. In each phase we add a new column to the construction, one by one, adding the $j$th column in the phase $j$. Let $x_{\ell}$ be the timestamp at the end of phase $j$. We will preserve the following invariant: It holds that $\operatorname{lab}_{\ell}(r)=\ell_{\operatorname{root}}$, $\operatorname{lab}_{\ell}(r_{\boldsymbol{B}})=\ell_{\operatorname{black}}$, $\operatorname{lab}_{\ell}(g_{1})=\ell_{\operatorname{g_{1}}}$, $\operatorname{lab}_{\ell}(g_{2})=\ell_{\operatorname{g_{2}}}$, and
$\operatorname{lab}_{\ell}(v^{4}(X^{i}_{j}))=\operatorname{lab}_{\ell}(v^{5}(X^{i}_{j}))=\operatorname{lab}_{\ell}(v^{6}(X^{i}_{j}))=i$
for all $i\in[n^{\prime}]$. All other vertices in $V_{\ell}$ are labeled $\ell_{\operatorname{forget}}$ at $\ell$.
@@
Finally, We are left with the vertices of path gadgets, together with the vertex $r_{\boldsymbol{B}}$. We show that we can define a proper 2-coloring $\gamma$ of these vertices that extends the coloring $\{r_{\boldsymbol{B}}\mapsto\boldsymbol{B}\}$. First, we assign the color $\boldsymbol{B}$ to all vertices that were colored black, and the color $\boldsymbol{W}$ to all vertices colored white in the construction of $G$. Since $r_{\boldsymbol{B}}$ separates the different path sequences in $G^{\prime}$, it suffices to fix $i\in[n^{\prime}]$ and show that we can color all vertices of the path gadgets on the $i$th path sequence consistently with the fixed coloring of $r_{\boldsymbol{B}}$. We color all path gadgets $X_{i}^{j}$ in the same way for all values of $j$. Recall that $S$ includes the same vertices in all path gadgets on the $i$th path sequence. Assume that $X_{i}^{j}$ is the $h$th path gadget on the $\ell$th bundle, and let $\pi_{\ell}(h)=s=(\chi_{1},\chi_{2},\chi_{3})$. Let $\operatorname{next}(s)=(\chi_{4},\chi_{5},\chi_{6})$. Fix $q\in[6]$, let $\chi=\chi_{q}$. We show how to color the simple path gadget $Y=Y_{q}$. If $\chi\in\{\boldsymbol{B},\boldsymbol{W}\}$, let $\chi^{\prime}=\{\boldsymbol{B},\boldsymbol{W}\}\setminus\chi$. We assign the color $\chi$ to $v$. The vertex $w$ is isolated, and hence, can be colored arbitrarily. Finally, we assign the color $\chi^{\prime}$ to the two vertices created to subdivide the edges $\{v,x_{\boldsymbol{B}}\}$ and $\{v,x_{\boldsymbol{W}}\}$. On the other hand, if $\chi\in\{\boldsymbol{C},\boldsymbol{D}\}$, then the vertices $u,z$ and the vertices created to subdivide edges are all isolated, and can be colored arbitrary.
@@
We show that $|S_{Q}|\geq{b}(Q)$ must hold. The equality, and the fact that $S_{Q}$ does not contain any other vertices follows from the tightness of ${b}$, since all components $Q$ are pairwise disjoint.
@@
The set $S$ cannot contain any triangle vertices.
@@
Let $X$ be some path gadget. Let $s=(\chi_{1},\chi_{2},\chi_{3})$ be the state of $X$, and let $\operatorname{next}(s)=(\chi_{4},\chi_{5},\chi_{6})$. Then, it must holds that $\chi_{i}$ is the state of $Y_{i}$ for all $i\in[6]$.
@@
Assume that the Connected Odd Cycle Transversal problem can be solved in time $\mathcal{O}^{*}((12-\varepsilon)^{k})$ given a $k$-expression $\mu$ of $G$. We show that there exists a positive constant $\delta$, such that the $d$-SAT problem can be solved in time $\mathcal{O}^{*}((2-\delta)^{n})$ for all values of $d$, which contradicts SETH. Given an instance $I$ of the $d$-SAT problem, first we build the graph $G$, described above and then we run the given algorithm on the instance $(G,{b})$. It holds from Lemma 6.2 and Lemma 6.10 that $G$ admits a connected odd cycle transversal of size ${b}$ if and only if $I$ is satisfiable.
@@
Let $\alpha=\log_{12}(12-\varepsilon)$. We choose $t_{0}$ to be any constant integer larger than $\frac{\alpha}{\log_{12}2\cdot(1-\alpha)}$ Then it holds that

### 26 Feb 2024 v2

@@ 5.4 Isolation lemma
Assume that $G$ contains a connected odd cycle transversal of size $\overline{b}$. Fix $W=8\cdot|\mathcal{V}^{CJ}|$, and let $\omega$ is chosen uniformly at random in $[W]^{\big{(}\mathcal{V}^{CJ}\times[4]\big{)}}$. Let $\tau$ be a minimum weight valid action-sequence of cost $\overline{b}$ generating a pair $([0],c)$ for any coloring $c$. Then $\tau$ is unique with probability at least one half.
@@
Let $P$ be the family of all valid action-sequences of cost $\overline{b}$ generating a pair $([0],c)$ for any coloring $c$. For each $\tau\in P$, the weight of $\tau$ is given by $\sum\limits_{x\in\mathcal{V}^{CJ}}\omega(x,\tau(x))$. Let $U=\mathcal{V}^{CJ}\times[4]$, and for $\tau\in P$, let $F_{\tau}=\{(x,\tau(x))\colon x\in\mathcal{V}^{CJ}\}$. Let =$\mathcal{F}=\{F_{\tau}\colon\tau\in P\}$. Then $\tau\mapsto F_{\tau}$ is a bijection from $P$ to $\mathcal{F}$. Moreover, it holds that $\omega(\tau)=\omega(F_{\tau})$ for all $\tau\in P$. Finally, by Lemma 5.24, it holds that $\omega$ isolates $\mathcal{F}$ with probability
@@ 5.5 Correctness of the algorithm
Let $W=8\cdot|\mathcal{V}^{CJ}|$. First, choose a weight function $\omega:\mathcal{V}^{CJ}\times[4]\rightarrow W$, by choosing $\omega(x,\ell)\in[W]$ independently and uniformly at random for each $(x,\ell)\in\mathcal{V}^{CJ}\times[4]$. For each vertex $v\in V$ the algorithm repeats the following: the algorithm chooses $v_{0}=v$, and then computes all tables $T{\big{[}{x,{b},{w}}\big{]}}$ for all $x\in\mathcal{V}$, and all values ${b}\in[\overline{b}]$ and ${w}\in[4\cdot|\mathcal{V}^{CJ}|\cdot W]$ in a bottom-up manner over $\mathcal{T}$. The algorithm returns yes, if for any choice of $v_{0}$ there exists a coloring $c\in\mathscr{C}$ and a weight ${w}$ such that $T{\big{[}{r,\overline{b},{w}}\big{]}}\big{[}{([0],c)}\big{]}=1$. Otherwise, it returns no.
@@
On the other hand, if a solution of size $\overline{b}$ exists. Let $v_{0}$ be any vertex in any solution of size $\overline{b}$. Then by Lemma 5.25 there exists a weight ${w}$ and a coloring $c$, such that with probability at least $1/2$ there exists a unique action-sequence at $r$ of cost $\overline{b}$ and weight ${w}$ generating the pair $([0],c)$, which implies by Corollary 5.22 that $T{\big{[}{r,\overline{b},{w}}\big{]}}\big{[}{([0],c)}\big{]}=1$. Hence, with probability at least $1/2$ the algorithm outputs yes. ∎
@@ 6.1 Construction of the graph
By adding the triangle $\Delta({u},{v})$ to a graph $G$ for two vertices $u,v\in V(G)$, we mean that we add the edge $\{u,v\}$, and we add a new vertex $w_{u,v}$ and make it adjacent to $u$ and $v$ only. When we say we add a triangle at a vertex $v$, we mean that we add two new vertices $u_{v},w_{v}$, and make them adjacent to each other and to $v$ only. Vertices introduced when adding triangles are called triangle vertices. For a set of vertices $S\subseteq V$, we denote by $\Delta{S}$ the set consisting of $S$ and all vertices added to create triangles at single vertices of $S$ or between pairs of vertices in $S$.
We build $G$ by combining copies of fixed components of constant sizes, called gadgets. We distinguish three kinds of gadgets in $G$, namely path gadgets, decoding gadgets, and clause gadgets. We distinguish a single vertex $r$ in $G$ called the root vertex, with a triangle added to it, and another single vertex $r_{\boldsymbol{B}}$. When we say that we color a vertex white in $G$, it means that we add an edge between this vertex and $r_{\boldsymbol{B}}$. When we say we color a vertex $u$ black, we mean that we introduce a new vertex $u^{\prime}$, we color it white, and we add an edge between $u$ and $u^{\prime}$. We also distinguish two vertices $g_{1}$ and $g_{2}$ in $G$, called guards. We add a triangle at each of them.
@@
Let $t=\left\lceil t_{0}\cdot\log_{12}(2)\right\rceil$ and $n^{\prime}=s\cdot t$. The graph $G$ consists of $n^{\prime}$ path-like structures, called path sequences. Each path sequence is a sequence of $c=(11\cdot n^{\prime}+1)\cdot m$ path gadgets. Consecutive path gadgets are connected by cuts of size $9$, given as bicliques between three vertices of one gadget, and three of the other, giving a path sequence its narrow path-like structure. The first path gadget of each path sequence is adjacent to $g_{1}$, while the last path gadget is adjacent to $g_{2}$.
Let $P_{1},\dots,P_{n^{\prime}}$ be the set of all path sequences in $G$. We group them into $s$ groups of size $t$ each, calling each of these groups a path bundle. Hence, the $i$th path bundle $\mathcal{B}_{i}$ consists of the path sequences $P_{(i-1)\cdot t+1}\dots P_{i\cdot t}$, and it corresponds to the variable group $V_{i}$. For each $i\in[s],j\in[c]$, we call the set of all $j$th path gadgets on all path sequences in the path bundle $\mathscr{B}_{i}$ a (simple) bundle $B_{i}^{j}$.
For a path sequence $P_{i}$, let $X_{i}^{1},\dots X_{i}^{c}$ be the path gadgets on the path sequence $P_{i}$ from left to right. We divide each path sequence $P_{i}$ into $(11\cdot n^{\prime}+1)$ groups of size $m$ each, called segments $S_{i}^{1},\dots,S_{i}^{11\cdot n^{\prime}+1}$, where $S_{i}^{j}=X_{i}^{(j-1)\cdot m+1},\dots,X_{i}^{j\cdot m}$. For $j\in[11\cdot n^{\prime}+1]$, we call the set consisting of the $j$th segment of each path sequences a section $\mathcal{S}_{j}$, i.e. $\mathcal{S}_{j}=\{S_{i}^{j}\colon i\in[n^{\prime}]\}$. Finally, for $j\in[c]$, we call the set consisting of the $j$-th path gadget of each path sequence a column $\operatorname{Col}_{j}=\{X_{i}^{j}\colon i\in[n^{\prime}]\}$. See Figure 1 for illustration.
@@
We attach a decoding gadget $D_{i}^{j}$ to each bundle $B_{i}^{j}$ for $i\in[s]$ and $j\in[c]$, and we attach a clause gadget $Z_{j}$ to each column $\operatorname{Col}_{j}$ for $j\in[c]$. A decoding gadget is only adjacent (possibly through triangles) to vertices of the path gadgets in its corresponding bundle, to the clause gadget in its column, and to $r$. A clause gadget is only adjacent to vertices in the decoding gadgets attached to bundles in its column.
We define the set =$\mathscr{I}=\{\boldsymbol{C},\boldsymbol{D},\boldsymbol{B},\boldsymbol{W}\}$, and call it the set of simple states. We define the set of states $\mathcal{S}$ as the set of tuples $\mathcal{S}=\{s_{1},\dots s_{12}\}\subseteq\mathscr{I}^{3}$ given by the following list:
@@
$\big{|}\mathcal{S}^{B}\big{|}=12^{t}\geq 12^{t_{0}\cdot\log_{12}2}=2^{t_{0}}\geq\big{|}\{\mathbf{T},\mathbf{F}\}^{V_{i}}\big{|},$
we can fix an arbitrary injective mapping $\tau:\{\mathbf{T},\mathbf{F}\}^{V_{i}}\rightarrow\mathcal{S}^{[t]}$. Since both the set of the variables, and all path gadgets in each bundle are ordered (the former by index, and the latter by the index of the path sequence they belong to), we can assume that the same mapping is defined for all variable groups $V_{i}$, and all bundles $B_{i}^{j}$. Now we describe the different types of gadgets.
@@
A path gadget $X$ consists of six simple path gadgets $Y_{1},\dots,Y_{6}$. We denote a vertex $x$ in $Y_{i}$ by $x^{i}$. Hence, by $v^{1}$ we denote the vertex $v$ in $Y_{1}$, and by $x_{\boldsymbol{C}}^{3}$ we denote the vertex $x_{\boldsymbol{C}}$ in $Y_{3}$. We call the vertices $v^{1},v^{2},v^{3}$ the entry vertices, and the vertices $v^{4},v^{5},v^{6}$ the exit vertices of $X$. These vertices play a special role in $G$, since they connect consecutive path gadgets on a path sequence. In particular, we add a biclique between the exit vertices of each path gadget (except for the last one), and the entry vertices of the following path gadget. The entry vertices of the first path gadget of each path sequence are all adjacent to $g_{1}$, and the exit vertices of the last path gadget are all adjacent to $g_{2}$.
For each $s\in\mathcal{S}$, we add a vertex $z_{s}$ to $X$ and make it adjacent to $r$. We call the resulting vertices the transition vertices of $X$. We add all triangles $\Delta({z_{s}},{z_{s^{\prime}}})$ for all $s\neq s^{\prime}\in\mathcal{S}$. Finally, for each $s\in\mathcal{S}$, let $s=(\chi_{1},\chi_{2},\chi_{3})$ and $\operatorname{next}(s)=(\chi_{4},\chi_{5},\chi_{6})$, where $\chi_{i}\in\mathscr{I}$ for all $i\in[6]$. For each $i\in[6]$ we add all triangles $\Delta({z_{s}},{x_{\chi}^{i}})$, for all $\chi\neq\chi_{i}\in\mathscr{I}$. That means we add a triangle between a transition vertex $z_{s}$ and a state vertex $x_{\chi}^{i}$ in $Y_{i}$, if $\chi_{i}\neq\chi$. This completes the description of a path gadget. Given a vertex $v$ in a path gadget $X$, if not clear from the context, we write $v(X)$ to indicate the path gadget it belongs to. In total, a path gadget contains $6\cdot 10+12=72$ non-triangle vertices.
@@
Let $D=D_{i}^{j}$, and let $X_{1},\dots X_{t}$ be the path gadgets on the bundle $B_{i}^{j}$. For each mapping $\pi\in\Pi$ and for each $i\in[t]$ and $s\in\mathcal{S}$, we add the triangle $\Delta({z_{s}(X_{i})},{u_{\pi}})$ if it holds that $\pi(i)\neq s$. Given a vertex $v$ in a decoding gadget $D$, if not clear from the context, we write $v(D)$ to indicate the decoding gadget it belongs to. In total, a decoding gadget contains $2\cdot 12^{t}$ non-triangle vertices.
@@
Now we define the target value ${b}$ in the resulting instance $(G,{b})$ of the Connected Odd Cycle Transversal problem. First, we define a family of disjoint components $\mathcal{F}$ in $G$ (mostly the gadgets of $G$). For each component $X\in\mathcal{F}$, we will define a family of disjoint subsets of $X$ (denoted $\beta(X)$). For each set $S\in\beta(X)$, we will define a lower bound ${b}(S)$ on the size of a connected odd cycle transversal in $G[S]$. We define ${b}(X)$ as the sum of ${b}(S)$ for all $S\in\beta(X)$. Finally, we define ${b}$ as the sum of ${b}(X)$ for all $X\in\mathcal{F}$.
@@ 6.2 Proof of correctness
Now we are ready to describe the linear $k$-expression of $G$. First, we introduce the vertices $r,r_{\boldsymbol{B}},g_{1}$ and $g_{2}$ with labels $\ell_{\operatorname{root}}=n^{\prime}+3$, $\ell_{\operatorname{black}}=n^{\prime}+4$, $\ell_{\operatorname{g_{1}}}=n^{\prime}+5$, and $\ell_{\operatorname{g_{2}}}=n^{\prime}+6$ respectively. These are the only vertices that will be assigned the corresponding labels at any timestamp, and will never be relabeled. Hence, for a vertex $u$ of a unique label at any timestamp, we can make it adjacent to any of them through a join operation without changing the adjacency of any other vertex of the graph. We introduce the private neighbors of $r,g_{1},$ and $g_{2}$, using a different label for each introduced vertex. We create the corresponding triangles using join operations, and then we forget all these private neighbors. Finally, we reserve the label $\ell_{\operatorname{white}}=n^{\prime}+7$ for creating white vertices when coloring some vertices black. For a vertex $v$ of a unique label $\ell$, when we say we color $v$ white we mean that we join the labels $\ell$ and $\ell_{\operatorname{black}}$, while by coloring $v$ black, we mean that we introduce a vertex $w$, labeled $\ell_{\operatorname{white}}$, and then apply two join operations: one between $\ell_{\operatorname{white}}$ and $\ell_{\operatorname{black}}$, and another between $\ell_{\operatorname{white}}$ and $\ell$. After that we forget the label $\ell_{\operatorname{white}}$.
We divide the rest of the construction into phases. In each phase we add a new column to the construction, one by one, adding the $j$th column in the phase $j$. Let $x_{h}$ be the timestamp at the end of phase $j$. We will preserve the following invariant: It holds that $\operatorname{lab}_{h}(r)=\ell_{\operatorname{root}}$, $\operatorname{lab}_{h}(r_{\boldsymbol{B}})=\ell_{\operatorname{black}}$, $\operatorname{lab}_{h}(g_{1})=\ell_{\operatorname{g_{1}}}$, $\operatorname{lab}_{h}(g_{2})=\ell_{\operatorname{g_{2}}}$, and
$\operatorname{lab}_{h}(v^{4}(X_{i}^{j}))=\operatorname{lab}_{h}(v^{5}(X_{i}^{j}))=\operatorname{lab}_{h}(v^{6}(X_{i}^{j}))=i$
for all $i\in[n^{\prime}]$. All other vertices in $V_{h}$ are labeled $\ell_{\operatorname{forget}}$ at $x_{h}$.
@@
Finally, We are left with the vertices of path gadgets, together with the vertex $r_{\boldsymbol{B}}$. We show that we can define a proper 2-coloring $\gamma$ of these vertices that extends the coloring $\{r_{\boldsymbol{B}}\mapsto\boldsymbol{B}\}$. First, we assign the color $\boldsymbol{B}$ to all vertices that were colored black, and the color $\boldsymbol{W}$ to all vertices colored white in the construction of $G$. Since $r_{\boldsymbol{B}}$ separates the different path sequences in $G^{\prime}$, it suffices to fix $i\in[n^{\prime}]$ and show that we can color all vertices of the path gadgets on the $i$th path sequence consistently with the fixed coloring of $r_{\boldsymbol{B}}$. We color all path gadgets $X_{i}^{j}$ in the same way for all values of $j$. Recall that $S$ includes the same vertices in all path gadgets on the $i$th path sequence. Assume that $X_{i}^{j}$ is the $h$th path gadget on the $\ell$th bundle, and let $\pi_{\ell}(h)=s=(\chi_{1},\chi_{2},\chi_{3})$. Let $\operatorname{next}(s)=(\chi_{4},\chi_{5},\chi_{6})$. Fix $q\in[6]$, let $\chi=\chi_{q}$. We show how to color the simple path gadget $Y=Y_{q}$. If $\chi\in\{\boldsymbol{B},\boldsymbol{W}\}$, let $\chi^{\prime}=\{\boldsymbol{B},\boldsymbol{W}\}\setminus\chi$. We assign the color $\chi$ to $v$. The vertex $w$ is isolated, and hence, can be colored arbitrarily. Finally, we assign the color $\chi^{\prime}$ to the two vertices added to subdivide the edges $\{v,x_{\boldsymbol{B}}\}$ and $\{v,x_{\boldsymbol{W}}\}$. On the other hand, if $\chi\in\{\boldsymbol{C},\boldsymbol{D}\}$, then the vertices $u,z$ and the vertices added to subdivide edges are all isolated, and can be colored arbitrary.
@@
We show that $|S_{Q}|\geq{b}(Q)$ must hold. The equality and the fact that $S_{Q}$ does not contain any other vertices follows from the tightness of ${b}$, since all components $Q$ are pairwise disjoint.
@@
Any solution $S$ of size ${b}$ cannot contain any triangle vertices.
@@
Let $X$ be some path gadget. Let $s=(\chi_{1},\chi_{2},\chi_{3})$ be the state of $X$, and let $\operatorname{next}(s)=(\chi_{4},\chi_{5},\chi_{6})$. Then, it must hold that $\chi_{i}$ is the state of $Y_{i}$ for all $i\in[6]$.
@@
Assume that the Connected Odd Cycle Transversal problem can be solved in time $\mathcal{O}^{*}((12-\varepsilon)^{k})$ given a $k$-expression $\mu$ of $G$. We show that there exists a positive constant $\delta$, such that the $d$-SAT problem can be solved in time $\mathcal{O}^{*}((2-\delta)^{n})$ for all values of $d$, which contradicts SETH. Given an instance $I$ of the $d$-SAT problem, first we build the graph $G$, described above, and then we run the given algorithm on the instance $(G,{b})$. It holds from Lemma 6.2 and Lemma 6.10 that $G$ admits a connected odd cycle transversal of size ${b}$ if and only if $I$ is satisfiable.
@@
Let $\alpha=\log_{12}(12-\varepsilon)$. We choose $t_{0}$ to be any constant integer larger than $\frac{\alpha}{\log_{12}2\cdot(1-\alpha)}$. Then it holds that